OpenStudy (patrickweewoos):

Which value of x makes the following equation true? 5.7(x - 2) = -0.3x + 30.6 9 6 7 10

OpenStudy (patrickweewoos):

@iGreen

OpenStudy (patrickweewoos):

@puppylover234

OpenStudy (anonymous):

the answer is 7

OpenStudy (patrickweewoos):

Which value of x makes the following equation true? 3(x - 4) - 16 = 2(x - 8) + 6 10 18 14 26

OpenStudy (anonymous):

5.7x+0.3x=30.6+11.4 6x=42 x=7 its the perfect answer

OpenStudy (shinalcantara):

first thing to do is to distribute 5.7 to (x-2) then you'll have 5.7(x) - (5.7)(2) = -0.3x + 30.6 5.7x - 11.4 = -0.3x + 30.6 5.7x + 0.3x = 30.6 + 11.4 6x = 42 x = 7

OpenStudy (shinalcantara):

3(x - 4) - 16 = 2(x - 8) + 6 3x - 12 - 16 = 2x - 16 + 6 3x -2x = -16 + 6 + 16 + 12 x = 18

OpenStudy (patrickweewoos):

create a linear equation with a solution of x = 10.

OpenStudy (patrickweewoos):

@shinalcantara

OpenStudy (patrickweewoos):

Help?

OpenStudy (patrickweewoos):

create a linear equation with a solution of x = 10.

OpenStudy (anonymous):

cant help you with that one....got any others?

OpenStudy (patrickweewoos):

No i have to get passed this to get to others

OpenStudy (patrickweewoos):

Shinal do you know

OpenStudy (anonymous):

no clue :/

OpenStudy (shinalcantara):

4(3x+1) = 5(2x+5) - 1

OpenStudy (patrickweewoos):

@iGreen @lizgarza1998 @AnswerMyQuestions

OpenStudy (patrickweewoos):

Shinal can you make a linear equation it has a fraction

OpenStudy (anonymous):

@shinalcantara thanks for the help. Any more @PatrickWeeWoos ?

OpenStudy (patrickweewoos):

i have to get passed this to get others

OpenStudy (anonymous):

x=2.5

OpenStudy (patrickweewoos):

the numbers are 5,4,x,2 to put in a linear equation w/ a fraction for this

OpenStudy (patrickweewoos):

help